Molality

Calculate the molality of a solution that is prepared by mixing 25.5 mL of CH3OH (d= 0.792
g/mL) and 387 mL of CH3CH2CH2OH (d= 0.811 g/mL).
Answer #1
first decide which one is solvent.
since the amount of CH3CH2CH2OH is more than that of CH3OH,
CH3OH is the solute, while CH3CH2CH2OH is the solvent.
mass of CH3OH = density x volume = 25.5 x 0.792 = 20.2 g
moles of CH3OH = mass/ molar mass = 20.2/32 = 0.63 mol
mass of CH3CH2CH2OH = density x volume = 0.811 x 387 = 313.9 g = 0.3139 kg
molality = moles of CH3OH/mass of CH3CH2CH2OH = 0.63/0.3139 = 2.0 mol/kg


answered by: blah